# VaLID: Verification as Late Integration of Detections for LiDAR-Camera Fusion

### v1

@@ Abstract
Vehicle object detection is possible using both LiDAR and camera data. Methods using LiDAR generally outperform those using cameras only. The highest accuracy methods utilize both of these modalities through data fusion. In our study, we propose a model-independent late fusion method, VaLID, which validates whether each predicted bounding box is acceptable or not. Our method verifies the higher-performing, yet overly optimistic LiDAR model detections using camera detections that are obtained from either specially trained, general, or open-vocabulary models. VaLID uses a simple multi-layer perceptron trained with a high recall bias to reduce the false predictions made by the LiDAR detector, while still preserving the true ones. Evaluating with multiple combinations of LiDAR and camera detectors on the KITTI dataset, we reduce false positives by an average of 63.9%, thus outperforming the individual detectors on 2D average precision (2DAP). Our approach is model-agnostic and demonstrates state-of-the-art competitive performance even when using generic camera detectors that were not trained specifically for this dataset.
@@ I INTRODUCTION
We find that the VaLID method removes an average 63.9% of false positives, consistently improving overall detection average precision. Importantly, improvement is as effective using the general camera models as it is with the specifically trained method. In both cases results are competitive with state of the art.
The contribution of this paper is a model-independent late fusion method that combines LiDAR and camera detections.
Figure 1: We conduct our study on the KITTI dataset. The figure rows show (a) official 2D ground truth (b) detections from the specialized LiDAR model PVRCNN, (c) specialized camera model MonoDETR, and (d) open Vocabulary model GroundingDino. LiDAR generally produces too many false positives, seen most obviously in Column [i]. Camera models can help verify and reject false positives despite imperfections in their own detections.
@@ II-A Detection Using Camera Images
Cameras images are fundamentally a 2D data source, and 2D object detection is a well-studied area with many surveys and general purpose datasets [52, 21, 38, 16, 17, 28, 30, 15, 6]. Higher accuracy is possible in specific domains such as vehicle detection by designing specialized object detection methods which are trained on vehicle specific datasets [8, 33, 3]. Since autonomous driving requires 3D awareness, there have also been attempts to detect 3D bounding boxes from camera images. However, accurately estimating depth without dedicated sensors is challenging, especially for distant objects. Early approaches like Deep3DBox leverage 2D bounding boxes and geometric constraints to infer 3D dimensions but struggle with long-range accuracy [23]. Similarly, MonoDETR improves detection by incorporating depth embeddings into a transformer model [48]. Other methods generate depth maps to assist with detection. For instance, MonoPGC refines pixel-wise depth estimation through cross-attention [44], and Pseudo-LiDAR simulates LiDAR input from depth maps to enhance detection performance [39]. While these techniques improve depth accuracy, they often introduce significant computational overhead, especially for complex or occluded scenes. Multi-task approaches like Deep MANTA predict both 2D and 3D attributes simultaneously, leveraging shared features to enhance efficiency, but they still face challenges in reliable depth estimation [5]. Unfortunately, camera-only methods, while effective in some scenarios, are limited by their inability to produce reliable depth estimation without additional sensors, and 3D bounding box accuracy remains below 2D bounding box accuracy. In our work, we use the higher accuracy 2D detections as one of the inputs to our late fusion method, which integrates data from multiple sensors.
@@ II-C Detection Using LiDAR-Camera Fusion
Figure 2: Precision-Recall curve (PRC) of our baseline single modality models on the KITTI moderate difficulty data set. Notice that both LiDAR models outperform all three of the camera models, and that the specialized camera detector outperforms the two general purpose camera-based object detectors.
@@
Figure 4: In our method called VaLID, we take LiDAR (L) boxes as the primary detections and use the camera (C) modality to verify whether each L box is an acceptable detection or not. The input vector is given as the bounding box dimensions of width $w$, height $h$, and center $(cx,cy)$, confidence scores $s$, and the measure of overlap $IoU$. We pass this vector through an MLP and output an accept or reject signal as a sigmoid value. The target loss is calculated considering the overlap of the ground truth with the given L box.
@@ III-B Model Detections Complement Each Other
Fig. 2 shows the overall performance of our baseline LiDAR and camera methods. Notice that both LiDAR methods outperform all three camera methods, and that the specialized camera model outperforms the two general models. Since the LiDAR methods outperform the camera methods, we take them as our primary detectors, and use the camera methods to augment their detection ability.
@@ IV-A Method Overview
We use camera model detections and a neural network to verify whether a particular LiDAR bounding box detection is valid. Fig. 4 provides an overview of our method. For each LiDAR detection, we look for a corresponding camera detection. If a matching camera box is found, the camera bounding box, the Intersection over Union (IoU) between the LiDAR and camera boxes, and the confidence scores of both detections are fed into a neural network, along with the LiDAR box. If no matching camera box is found, we input a zero for both the bounding box and the confidence score for the camera model.
Our goal is to determine whether to keep or discard each LiDAR (L) detection. We generate ground truth by calculating the IoU between the LiDAR box and the official KITTI ground truth, and follow the datasets standard by setting an IoU threshold of 0.7. Detections above this threshold are marked as valid (positive class), while those below are ignored. The inputs given would be:
$I=\{\text{Bbox}_{L},\text{Bbox}_{C},S_{L},S_{C},\text{IoU}_{(L,C)},[\text{IoU}_{(L,GD)}]\}$
@@
The entire input vector is then fed to a two-layer multi-layer perceptron (MLP) and trained with a high recall to provide a bias towards avoiding accidentally removing true positive bounding boxes while reducing false positives. This MLP is simple and architecturally identical in all of our experiments. However the activation weights need to be trained for each pair of LiDAR and camera methods since each baseline method has its own distributions of true and false positives and the goal of the MLP is to make specific tradeoffs in these distributions.
In general we desire over production of bounding boxes in the verification method as opposed to under production. Since the detected boxes of the open vocabulary model are not as precise as the specialized LiDAR and camera models, especially around occluded objects, we use a lower IoU threshold to allow more boxes to match.
@@ IV-B Implementation
To implement our late-fusion method, we combine two specialized LiDAR detectors (PVRCNN and TED) with various camera models: a specialized model (MonoDETR), a general-purpose model (YOLO-NAS), and an open-vocabulary model (GroundingDINO). The experiments are conducted on the KITTI dataset. We remove dataset images for which ground truth is not available, as well as those which have already been used to train our baseline models. This leaves 3769 images available for our experiments comprising the subset that the dataset providers call validation. We split these images into training and testing subsets. For PVRCNN, these images generate a training set of 12,303 LiDAR bounding box samples and test set of 12,817 bounding boxes. For TED-S, we obtain 9,016 training and 9,406 testing bounding boxes.
The final inputs to our model include details such as LiDAR bounding box dimensions, confidence scores, and IoUs from the camera methods. These inputs pass through a two-layer multi-layer perceptron (MLP) with ReLU activation function and a last sigmoid layer. This output decides whether to keep each LiDAR box or discard the detection. The training loss function compares these outputs with the ground truth labels. Since our focus is on reducing false positives without significantly disturbing the true positive count, we use class weights of 1:10 for negative and positive classes to maintain high recall. We train our models for 50 epochs with a binary cross entropy (BCE) loss and an Adam learning rate of 0.0001.
@@ V Results
We present the results of our methods in two parts: First, we show the impact of our fusion method on reducing false positives generated by LiDAR-only models. Second, we provide an overall comparison of the 2D Average Precision (AP), comparing our fusion method with single-modality models as well as state-of-the-art fusion methods.
@@ V-A False Positives Reduction
Table I shows the performance of LiDAR-only models in terms of true positive and false positive bounding box predictions. Note that both LiDAR models produce a large number of false positives. Our late fusion method is evaluated on each LiDAR model, using each of our three baseline camera detectors, as well as using combinations of camera detectors. Our method consistently and substantially reduces false positives across all combinations. This is made especially clear through the visualization in Figure 5. Our late fusion method achieves an average false positive reduction of 73.6% on PVRCNN and 54.2% on TED. Importantly, the true positive (TP) numbers have only a small reduction, indicating that the significant decrease in false positives does not come at the cost of many missed detections.
2D AP (%)
@@
MonoDETR[48] C Specialized 95.12 87.36 79.94
YOLO-NAS[29] C General 75.32 68.69 55.11
GroundingDino[16] C Open-Vocab 84.73 78.68 62.82
PVRCNN[31] L Specialized 98.30 94.44 94.08
Ours (PVRCNN+MonoDETR) LC Late-Fusion 98.88 95.11 92.89
Ours (PVRCNN+YOLO-NAS) LC Late-Fusion 98.88 95.14 92.91
Ours (PVRCNN+GDino) LC Late-Fusion 98.88 95.14 92.92
Ours (PVRCNN+MonoDETR+GDino) LCC Late-Fusion 98.90 95.20 92.96
Ours (PVRCNN+YOLO-NAS+GDino) LCC Late-Fusion 98.92 95.23 92.95
TED-S[41] L Specialized 98.88 96.73 94.96
Ours (TED+MonoDETR) LC Late-Fusion 99.14 93.34 93.35
Ours (TED+YOLO-NAS) LC Late-Fusion 99.15 93.35 95.41
Ours (TED+GDino) LC Late-Fusion 99.11 95.35 95.44
Ours (TED+MonoDETR+GDino) LCC Late-Fusion 99.18 95.46 95.51
Ours (TED+YOLO-NAS+GDino) LCC Late-Fusion 99.13 95.34 95.39
CLOCs [24] LC Late-Fusion 99.33 93.75 92.89
CMKD[9] LC Knowledge Distil. 96.08 92.43 87.11
LogoNet[13] LC Deep Fusion 97.04 89.44 89.20
TED-M[41] LC Deep Fusion 99.48 96.25 93.72
TABLE II: Perfomance on the Average Precision (AP) metric for our method on all difficulty sets on KITTI-val. Our late fusion method consistently outperforms the original specialized LiDAR and camera methods on the easy set, along with moderate on PVRCNN and hard on TED. Our method is also comparable to the SOTA deep fusion methods, despite being simpler and lightweight, without requiring special integrated training.
Figure 6: PRC curves showing the LiDAR method TED before and after adding validation with our method. Especially for the lower confidence predictions needed to produce high recall rates, the LiDAR method produces many false positives. Zoomed-in plots reveal our method’s effectiveness in reducing these false positives, increasing the average precision (AP).
@@ V-B Average Precision
Table II presents the 2D Average Precision (AP) across the Easy, Moderate, and Hard KITTI sets, comparing individual LiDAR models, camera models, and their fusion combinations. Our late fusion method shows improved performance compared to any single-modality models, outperforming PV-RCNN on the Easy and Moderate sets, and TED-S on the Easy and Hard sets. The general and open-vocabulary camera models YOLO-NAS and GroundingDINO, perform as well as the special purpose camera model MonoDETR. These models, not specifically trained on KITTI, achieve comparable performance when fused with LiDAR models, demonstrating the adaptability of our method across different types of detectors.
In comparison with several state-of-the-art (SOTA) fusion methods, our method shows comparable results. Some of the comparison methods show better performance for the Easy and Moderate sets, while our method demonstrates better results on the Hard set. Our late-fusion method can match more complex fusion methods in terms of accuracy while offering better adaptability and lower computational overhead.
To better understand our verification based method, Fig. 6 shows the Precision-Recall (PR) curves comparing TED-S with two of our fusion approaches. The PR curves show that at high recall values our method is able to increase precision by removing false positives.

### v2

@@ Abstract
Vehicle object detection benefits from both LiDAR and camera data, with LiDAR offering superior performance in many scenarios. Fusion of these modalities further enhances accuracy, but existing methods often introduce complexity or dataset-specific dependencies. In our study, we propose a model-adaptive late-fusion method, VaLID, which validates whether each predicted bounding box is acceptable or not. Our method verifies the higher-performing, yet overly optimistic LiDAR model detections using camera detections that are obtained from either specially trained, general, or open-vocabulary models. VaLID uses a simple multi-layer perceptron trained with a high recall bias to reduce the false predictions made by the LiDAR detector, while still preserving the true ones. Evaluating with multiple combinations of LiDAR and camera detectors on the KITTI dataset, we reduce false positives by an average of 63.9%, thus outperforming the individual detectors on 3D average precision (3DAP). Our approach is model-adaptive and demonstrates state-of-the-art competitive performance even when using generic camera detectors that were not trained specifically for this dataset.
@@ I INTRODUCTION
Figure 1: We conduct our study on the KITTI dataset. The figure rows show (a) official 2D ground truth (b) detections from the specialized LiDAR model PVRCNN, (c) specialized camera model MonoDETR, and (d) open Vocabulary model GroundingDino. LiDAR generally produces too many false positives, seen most obviously in Column [i]. Camera models can help verify and reject false positives despite imperfections in their own detections (e.g. Missed detection of partial car in Row[c]Col[ii], and Imperfect dimensions in Row[d]Col[iii]).
@@
We find that the VaLID method removes an average 63.9% of false positives, consistently improving overall detection average precision. Importantly, improvement is as effective using the general camera models as it is with the specifically trained method. In both cases, results are competitive with the state of the art.
Thus, the primary contributions of our work are:
  • We propose VaLID, a lightweight, model-adaptive late-fusion approach that focuses on reducing false positives by cross-validating them with camera detections.
  • Our method works with both specialized and general-purpose camera models without requiring individual model retraining or dataset-specific fine-tuning, making it adaptable to various detection backbones.
  • Our method uses a multi-layered perceptron (MLP) with confidence rescoring that balances simplicity and effectiveness, avoiding the complexity of deep-fusion methods while achieving competitive performance.
@@ II-A Detection Using Camera Images
Cameras images are fundamentally a 2D data source, and 2D object detection is a well-studied area with many surveys and general purpose datasets [53, 21, 39, 16, 17, 29, 31, 15, 6]. Higher accuracy is possible in specific domains such as vehicle detection by designing specialized object detection methods which are trained on vehicle specific datasets [8, 34, 3]. Since autonomous driving requires 3D awareness, there have also been attempts to detect 3D bounding boxes from camera images. However, accurately estimating depth without dedicated sensors is challenging, especially for distant objects. Early approaches like Deep3DBox leverage 2D bounding boxes and geometric constraints to infer 3D dimensions but struggle with long-range accuracy [23]. Similarly, MonoDETR improves detection by incorporating depth embeddings into a transformer model [49]. Other methods generate depth maps to assist with detection. For instance, MonoPGC refines pixel-wise depth estimation through cross-attention [45], and Pseudo-LiDAR simulates LiDAR input from depth maps to enhance detection performance [40]. While these techniques improve depth accuracy, they often introduce significant computational overhead, especially for complex or occluded scenes. Multi-task approaches like Deep MANTA predict both 2D and 3D attributes simultaneously, leveraging shared features to enhance efficiency, but they still face challenges in reliable depth estimation [5]. Unfortunately, camera-only methods, while effective in some scenarios, are limited by their inability to produce reliable depth estimation without additional sensors, and 3D bounding box accuracy remains below 2D bounding box accuracy. In our work, we use the higher accuracy 2D detections from camera images as one of the inputs to our late fusion method, which integrates data from multiple sensors.
Figure 2: Precision-Recall curve (PRC) of our baseline single modality models on the KITTI moderate difficulty data set. Notice that both LiDAR models outperform all three of the camera models and that the specialized camera detector outperforms the two general-purpose camera-based object detectors.
@@ II-C Detection Using LiDAR-Camera Fusion
Figure 4: In our method called VaLID, we take LiDAR (L) boxes as the primary detections and use the camera (C) modality to verify whether each L box is an acceptable detection or not. The input vector is given as the bounding box dimensions of width $w$, height $h$, and center $(cx,cy)$, confidence scores $s$, and the measure of overlap $IoU$. We pass this vector through an MLP and output an accept or reject signal as a sigmoid value. The target loss is calculated considering the overlap of the ground truth with the given L box. During the inference, the accepted box confidence is rescored by multiplying it with the output sigmoid probability in order to reduce the confidence of the lingering false positives.
@@ III-B Model Detections Complement Each Other
Fig. 2 shows the overall performance of our baseline LiDAR and camera methods. Notice that both LiDAR methods outperform all three camera methods, and that the specialized camera model outperforms the two general models. Since the LiDAR methods outperform the camera methods, we take them as our primary detectors and use the camera methods to augment their detection ability.
@@ IV-A Method Overview
We use camera model detections and a neural network to verify whether a particular LiDAR bounding box detection is valid. Fig. 4 provides an overview of our method. For each LiDAR detection, we look for a corresponding camera detection in which the overlap score (IoU) is 0.5 or higher. This threshold is intentionally set lower than the official KITTI benchmark of 0.7 to account for potential inaccuracies in the camera detection dimensions, ensuring that valid detections are not inadvertently discarded due to minor misalignments. If a matching camera box is found, the camera bounding box, the Intersection over Union (IoU) between the LiDAR and camera boxes, and the confidence scores of both detections are fed into a neural network, along with the LiDAR box. If no matching camera box is found, we input a zero for both the bounding box and the confidence score for the camera model. The inputs given would be:
$I=\{\text{Bbox}_{L},\text{Bbox}_{C},S_{L},S_{C},\text{IoU}_{(L,C)},[\text{IoU}_{(L,GD)}]\}$ (1)
@@
Our goal is to determine whether to keep or discard each LiDAR (L) detection. We generate ground truth by calculating the IoU between the LiDAR box and the official KITTI ground truth, and follow the dataset’s standard by setting an IoU threshold of 0.7 for the Car class. Detections above this threshold are marked as valid (positive class), while those below are ignored.
Validating LiDAR detections with camera detections could be approached with a simple thresholding heuristics. However, simple heuristics would fail to capture the complex, non-linear relationships between LiDAR and camera detections, such as bounding boxes with high confidence but low IoU or cases where one modality compensates for the other. A neural network allows this relationship to be learned from the data, rather than hand tweaked by the researcher for each new pair of detectors. Given the small, fixed-size feature vector in a structured format, an MLP is an efficient and effective choice. It can learn non-linear patterns and adapt to biases in false positives across different parametric bands, which are not the same in different detection models (Fig. 3). More complex models like CNNs or transformers, designed for high-dimensional inputs, would add unnecessary complexity without offering meaningful improvements.
@@
The input vector I (Eq. 1) is, therefore, fed to a two-layered MLP that is specifically designed to address the tradeoff between preserving true positives and reducing false positives. The MLP is trained with a high recall bias to minimize the risk of mistakenly discarding true positive detections while effectively filtering out false positives. This MLP is simple and architecturally identical in all of our experiments. However, the activation weights need to be trained for each pair of LiDAR and camera methods since each baseline method has its own distributions of true and false positives, and the goal of the MLP is to make specific tradeoffs in these distributions.
To ensure that enough verification boxes exist, we desire an over-production of bounding boxes in the camera-based verification methods as opposed to under-production, and thus include even boxes with low confidence. We observe that the open vocabulary camera model provides detected boxes that are not as precisely aligned as the specialized LiDAR and camera models, especially around occluded objects, and thus adjust the IoU threshold to ensure sufficient matching boxes exist.
Since we employ high-recall training, we expect some false positives will continue to be accepted by the MLP during inference. To mitigate their impact on the overall average precision (AP), we assign a final confidence ($S_{L}^{\prime}$)to each box by multiplying the accepted LiDAR box confidence score $S_{L}$ with the predicted output sigmoid probability $P_{S}$. This ensures that true positive boxes retain high confidence, while false positives with lower predicted probabilities are down-weighted, effectively balancing the AP.
3D AP / 2D AP (%)
Models Modality Model Type Easy Moderate Hard
MonoDETR[49] C Specialized 28.05 / 95.12 20.77 / 87.36 17.43 / 79.94
YOLO-NAS[30] C General N/A / 75.32 N/A / 68.69 N/A / 55.11
GroundingDino[16] C Open-Vocab N/A / 84.73 N/A / 78.68 N/A / 62.82
PVRCNN[32] L Specialized 91.94 / 98.30 82.97 / 94.44 82.44 / 94.08
Ours (PVRCNN+MonoDETR) LC Late-Fusion 91.96 / 98.91 83.47 / 95.45 82.89 / 93.11
Ours (PVRCNN+YOLO-NAS) LC Late-Fusion 92.30 / 99.09 83.83 / 95.83 82.91 / 93.33
Ours (PVRCNN+GDino) LC Late-Fusion 92.41 / 98.91 83.54 / 95.26 83.01 / 93.01
Ours (PVRCNN+MonoDETR+GDino) LCC Late-Fusion 92.37 / 99.15 83.91 / 95.87 83.08 / 93.37
Ours (PVRCNN+YOLO-NAS+GDino) LCC Late-Fusion 92.46 / 99.24 83.93 / 95.99 82.95 / 93.42
TED-S[42] L Specialized 95.23 / 98.88 87.98 / 96.73 86.28 / 94.95
Ours (TED+MonoDETR) LC Late-Fusion 95.40 / 99.23 88.69 / 95.64 86.77 / 95.59
Ours (TED+YOLO-NAS) LC Late-Fusion 95.18 / 99.12 88.73 / 95.79 86.64 / 95.66
Ours (TED+GDino) LC Late-Fusion 95.86 / 99.35 88.85 / 95.69 86.78 / 95.57
Ours (TED+MonoDETR+GDino) LCC Late-Fusion 95.59 / 99.40 88.94 / 95.93 86.80 / 95.74
Ours (TED+YOLO-NAS+GDino) LCC Late-Fusion 95.38 / 99.19 88.78 / 95.87 86.83 / 95.70
CLOCs [25] LC Late-Fusion 92.74 / 99.33 82.90 / 93.75 77.75 / 92.89
CMKD[9] C(+L) Knowledge Distil. 34.56 / 96.08 23.10 / 92.43 19.52 / 87.11
LogoNet[13] LC Deep Fusion 92.04 / 97.04 85.04 / 89.44 84.31 / 89.20
TABLE II: Perfomance on the Average Precision (AP) metric for our method on all difficulty sets on KITTI-val for the Car class. In general, the baseline camera methods do not produce 3D bounding boxes (N/A) or produce ones with very low performance (MonoDETR). Our late fusion method consistently outperforms the original specialized LiDAR and camera methods in 3DAP across all the difficulty levels. Our method is also comparable to the SOTA deep fusion methods, despite being simpler and lightweight, without requiring special integrated training on the individual models. The color coding in yellow emphasizes the degree of improvement on 3DAP (primary comparison metric) over the baselines.
@@ IV-B Implementation
To implement our late-fusion method, we combine two specialized LiDAR detectors (PVRCNN and TED) with various camera models: a specialized model (MonoDETR), a general-purpose model (YOLO-NAS), and an open-vocabulary model (GroundingDINO). The experiments are conducted on the KITTI dataset. We remove dataset images for which ground truth is not available, as well as those that have already been used to train our baseline models. This leaves 3769 images available for our experiments comprising the subset that the dataset providers call validation. We split these images into training and testing subsets. For PVRCNN, these images generate a training set of 12,303 LiDAR bounding box samples and test set of 12,817 bounding boxes. For TED-S, we obtain 9,016 training and 9,406 testing bounding boxes.
The final inputs to our model include details such as LiDAR bounding box dimensions, confidence scores, and IoUs from the camera methods. These inputs pass through a two-layer multi-layer perceptron with ReLU [24] activation function and a last sigmoid layer. This output decides whether to keep each LiDAR box or discard the detection. Since our focus is on reducing false positives without significantly disturbing the true positive count, we use class weights of 1:10 for negative and positive classes to maintain high recall. The model is optimized using a binary cross-entropy loss function, which compares the sigmoid outputs against the ground truth labels to predict the acceptance or rejection of a given LiDAR detection. We train our models for 50 epochs with an Adam learning rate of 0.0001. During the inference time, we rescore the accepted LiDAR box confidence score by multiplying it by the output sigmoid prediction probability, refining the final detection confidence.
@@ V Results
We present the results of our methods in two parts: First, we show the impact of our fusion method on reducing false positives generated by LiDAR-only models. Second, we provide an overall comparison of the Average Precision (3D/2DAP), comparing our fusion method with single-modality models as well as state-of-the-art fusion methods.
@@ V-A False Positives Reduction
Table I shows the performance of LiDAR-only models in terms of true positive and false positive bounding box predictions. Note that both LiDAR models produce a large number of false positives. Our late fusion method is evaluated on each LiDAR model, using each of our three baseline camera detectors, as well as using combinations of camera detectors. Our method consistently and substantially reduces false positives across all combinations. This is made especially clear through the visualization in Figure 5. Our late fusion method achieves an average false positive reduction of 73.6% on PVRCNN and 54.2% on TED. Importantly, the true positive (TP) numbers have only a small reduction, indicating that the significant decrease in false positives does not come at the cost of many missed detections. The results are qualitatively shown in Figure 6.
Figure 6: Qualitative results illustrating the reduction of false positives from our fusion method as compared to the specialized LiDAR PVRCNN. (a) denotes the ground truth bounding boxes in the KITTI-val set. In (b), red boxes represent the false positives generated by PVRCNN, while green boxes indicate the final accepted boxes from our late fusion method, demonstrating a significant decrease in false positives.
3D AP / 2D AP (%)
@@
PVRCNN[32] L Specialized 64.25 / 72.00 57.00 / 66.58 52.00 / 63.34
Ours (PVRCNN+YOLO-NAS) LC Late-Fusion 67.09 / 73.54 59.79 / 68.57 53.74 / 64.69
Ours (PVRCNN+GDino) LC Late-Fusion 65.64 / 72.02 58.42 / 67.30 52.82 / 63.42
Ours (PVRCNN+YOLO-NAS+GDino) LC Late-Fusion 67.28 / 73.97 60.36 / 68.01 54.31 / 63.58
TABLE III: Perfomance on the Average Precision (AP) metric for our method on all difficulty sets on KITTI-val for the Pedestrian class. Our late fusion method consistently outperforms the original specialized methods in 3DAP across all the difficulty levels. The color coding in yellow emphasizes the degree of improvement on 3DAP (primary comparison metric) over the baselines.
@@ V-B Average Precision
Table II presents the Average Precision (AP) for 3D object detection across the Easy, Moderate, and Hard KITTI sets on the Car class, comparing individual LiDAR models, camera models, and their fusion combinations. Our late fusion method shows consistently improved performance compared to any single-modality models, outperforming PV-RCNN and TED-S in 3D AP. The median improvement is +0.52 across all 3D-AP tests and difficulty levels, with an interquartile range of +0.43 to +0.86. While this improvement is modest, the performance on this dataset is already near saturation, leaving very little margin for error reduction. Hence, modest gains like this are typical of current state-of-the-art methods on KITTI for the Car class. To verify the generalizability of our method, we evaluate it on the Pedestrian class using models for which labels were available (Table III). Given a larger scope of error reduction for this class, we see that our fusion method demonstrates a median improvement of +2.31 with an interquartile range of +1.40 to +2.93 across all difficult levels.
Late fusion with general and open-vocabulary camera models (YOLO-NAS and GroundingDINO) perform as well as the tested special-purpose camera model (MonoDETR). This is surprising since the baseline AP of these detectors is much lower than the specially trained model. This provides evidence that general-purpose camera detectors can be used to enhance special-purpose LiDAR detectors.
Although our primary goal is to show that our late fusion method improves scores relative to the baseline methods, the table also provides numbers for comparison fusion methods. Our approach provides higher scores using a simpler method.